If f(x)= 3x-5 and g(x)= 2-x², find (f+g)(5)
Found 2 solutions by Theo, josgarithmetic:
Answer by Theo(13342)

f(x) = 3x-5
g(x) = 2-x^2
(f+g)(x) = 3x-5 + 2-x^2
simplify this and arrange in descending order of degree to get:
(f+g)(x) = -x^2+3x-3
(f+g)(5) = -5^2+3*5-3 = -25+15-3 = -13.

to see if this is true, solve for f(5) and g(5) separately and then add them together.

f(5) = 3*5-5 = 10
g(5) = 2-5^2 = -23
f(5) + g(5) = 10-23 = -13.

the value is the same.
(f+g)(5) give you the same value as f(5)+g(5).


Answer by josgarithmetic(39615)